Session: Oct 2011 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Sat Oct 29 23:32:44 2011) Success:
no known problems

Onsala (Michael Lindqvist at Sun Oct 30 18:13:24 2011) Minor failures:
Time-outs on the Mark5. Module(s) problems? See log for details.

Westerbork (Geert Kuper at Mon Oct 31 17:47:00 2011) Success:
RT7 in TADUMax.

Torun (Magda Kunert-Bajraszewska at Wed Nov 02 12:29:37 2011) Success:
BBC6 was unlocked during this experiment. No other known problems.

Yebes (Susana García-Espada at Fri Nov 04 10:08:34 2011) Success:
Not known problems.

Effelsberg (Uwe Bach at Mon Nov 07 11:45:58 2011) Minor failures: TSYS
16:58-17:00 failure elevation control, 2min lost 17:09-17:12 failure elevation control, 3min lost 17:10-21:00 noise diode didn't switch on. No calibration, but data is fine. Will introduce some reasonable numbers to the antabfs-file.

Shanghai (Bo Xia at Fri Nov 18 01:46:30 2011) Success: